Someone please help, I am totally confused with the following.
a) show the steps you would take to solve the following
4-(x/x-2)=(-2/x-2)
b) What potential solution did you obtained? Explain why this is not a solution.
Thank you in advance for you help

Answer by stanbon(57387)

a) show the steps you would take to solve the following
4-(x/x-2)=(-2/x-2)
Multiply thru by (x-2) to get:
4(x-2) - x = -2
4x-8 - x = -2
3x = 6
x = 2
---------
Checking:
4 - (2/0) = -2/0
Division by zero is not allowed; it is "undefined".
===================
Cheers,
Stan H.